OpenStudy (camerondoherty):

I think its B... The following table shows the values of f(x) and g(x) for different values of x: x 1 2 3 4 5 f(x) 2 4 8 16 32 g(x) 2 2 2 2 2 Which of the following best describes the function f(x) + g(x)? 2^(x + 2) 2^x + 2 2^x + 2x 2(2^x)

hartnn (hartnn):

can you find function f(x) in terms of x ? g(x) is clearly, just 2

OpenStudy (camerondoherty):

f(x)=x times 2

hartnn (hartnn):

f(1) = 2 = 2^1 f(2) = 4= 2^2 f(3) = 8 = 2^3 f(4) =16 = 2^4 ... do you get any hint from here ?

hartnn (hartnn):

if it was 2x, then f(3) would be 2*3 =6 and not 8

OpenStudy (camerondoherty):

its like 2*2 then 4*2 then 8*2 etc...

hartnn (hartnn):

yep, but in terms on 'x', it will be \(2^x\) and your answer B is correct :)

OpenStudy (camerondoherty):

Oh ok Thank You!

hartnn (hartnn):

welcome ^_^